OpenStudy (anonymous):

Will give medal! Need help with logs! find x when log (x^2 – 9) – 1 = log (x + 3)

OpenStudy (anonymous):

start with \[\log(x^2-9)-\log(x+3)=1\]

OpenStudy (anonymous):

then, using \[\log(A)-\log(B)=\log(\frac{A}{B})\] rewrite the left hand side as \[\log(\frac{x^2-9}{x+3})=1\]

OpenStudy (anonymous):

right, I got that far but I dont know how the answer is x=13 =/

OpenStudy (anonymous):

I got x=4

OpenStudy (anonymous):

did you get to \[\log(x-3)=1\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh i see, i guess you did but it is not \(x-3=1\) which would make \(x=4\) it is \[\log(x-3)=1\]

OpenStudy (anonymous):

rewrite in equivalent exponential form as \[x-3=10\]

OpenStudy (anonymous):

huh, where did that 10 come from?

OpenStudy (anonymous):

good question \[\log(\spadesuit)=\diamondsuit \iff 10^{\diamondsuit}=\spadesuit\]

OpenStudy (anonymous):

base of \(\log(whatever)\) is 10, i.e. it is \(\log_{10}(x-3)=1\) and so \(x-3=10^1\)

OpenStudy (anonymous):

like if you had \(\log(x-1)=2\) then \(x-1=10^2\) etc

OpenStudy (anonymous):

so it would be